Reports 1-1 of 1 Clear search Modify search
AdV-DAQ (Data Acquisition and Global Control)
masserot, viret - 14:20 Monday 17 November 2025 (68173) Print this report
DAQ Timing recovery

Due to a  thunderstorm, the Timing rack was switched off : no more Timing signals available for th ITF.  From the DAQ point of view the recovery was complete around 13H-UTC

 

All the TOLM devices were reconfigured : ADC7674, DBox and TOLM-PCIe mezzanine . 

To allows theses operations

  • all the loops were openned and all the Acl tasks were stopped .
  • The Acl servers have been restarted with theirs standard configurations
    • only the SIB2, SPRB, SDB2, SNEB and SWEB LC and SBE loops have been closed
    • for the EIB, LSC_Etalon_Acl, the  {CEB,NEB,WEB}_ALS_BPC none loops are closed.
    • for the TCS_CHROCC_PR the last DAC value has been restored , to be check by the expert
Comments to this report:
masserot - 8:34 Tuesday 18 November 2025 (68178) Print this report

The power supply for the Timing rack located in the DAQ was restored thanks to F.Berni and M.D'Andrea

Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×